OpenStudy (sleepyjess):

@Abhisar What would be the resulting molarity of a solution made by dissolving 45.7 grams of Mg(OH)2 in enough water to make a 1087-milliliter solution?

TheSmartOne (thesmartone):

Use the molar mass to find moles of Mg(OH)2. Convert the volume of the solution to Liters.

OpenStudy (sleepyjess):

.78 moles

OpenStudy (abhisar):

\(\sf \huge \boxed{Molarity = \frac{Moles~of~solute}{Volume~of~solution~in~litre}}\)

OpenStudy (sleepyjess):

that is just confusing me more... aahh :(

OpenStudy (sleepyjess):

so .78 would be moles of solute?

OpenStudy (abhisar):

First calculate the number of moles of 45.7 grams Mg(OH)2

OpenStudy (sleepyjess):

.78 moles

TheSmartOne (thesmartone):

:P

OpenStudy (abhisar):

one min

OpenStudy (sleepyjess):

I CAN'T WAIT A MINUTE! jk

OpenStudy (sleepyjess):

:D

TheSmartOne (thesmartone):

lol Abhi is timing you :P

OpenStudy (abhisar):

Yes, that's correct

OpenStudy (abhisar):

Now calculate the molarity by the formula i just told you.

OpenStudy (sleepyjess):

.00071757??

OpenStudy (abhisar):

1087 milliliter = how many litre?

OpenStudy (sleepyjess):

1.087

OpenStudy (sleepyjess):

.7175713

OpenStudy (abhisar):

So, Molarity = 0.78/1.087 = ?

OpenStudy (abhisar):

Jessie J, got it?

OpenStudy (sleepyjess):

lol yeah, I said it before you asked :)

OpenStudy (sleepyjess):

oh, can you check one other?

OpenStudy (abhisar):

Probably yes

OpenStudy (sleepyjess):

If 2.1 moles of CaCl2 are dissolved in enough water to make a 0.92-liter solution, what is the resulting molarity? I got approx 2.3

OpenStudy (abhisar):

This one so easy :P

OpenStudy (sleepyjess):

2.1/.92 right? :)

OpenStudy (abhisar):

2.1/0.92 =?

OpenStudy (sleepyjess):

LOL

OpenStudy (abhisar):

Yes \(\huge \checkmark\)

OpenStudy (abhisar):

\(\approx\)2.3

OpenStudy (sleepyjess):

:)

OpenStudy (abhisar):

:p :D

OpenStudy (abhisar):

Nom Nom Nom ^_^

OpenStudy (sleepyjess):

you are eating me? :o

OpenStudy (abhisar):

Yush.. Burp :D